OpenStudy (study_buddy99):

help me evaluate these 3 log problems?

OpenStudy (study_buddy99):

\[\log_\frac{ 2 }{ ?3 }\frac{ 27 }{ 8 }\]

OpenStudy (study_buddy99):

\[\log_x \frac{ 1 }{ 9 }\]

OpenStudy (study_buddy99):

\[\log_4x=-\frac{ 1 }{ 2 }\] *calculators are not allowed*

OpenStudy (study_buddy99):

the second question is also equal to -1/2

OpenStudy (jojokiw3):

First question?

OpenStudy (jojokiw3):

\[(\log_a x = y) \rightarrow (a^y = x)\] Try and use this to solve them.

OpenStudy (jojokiw3):

Logarithms are basically asking, to what power does "a" need to be raised to in order for you to get "x," thus the conversion equation above.

OpenStudy (study_buddy99):

I don't understand how to do that for the first question

OpenStudy (study_buddy99):

wait I think got it

OpenStudy (jojokiw3):

What's the answer for the second question? Just to make sure you get the concept?

OpenStudy (study_buddy99):

\[\frac{ 2 }{ 3 }^x=\frac{ 27 }{ 8 }\]

OpenStudy (study_buddy99):

would that be the start to number one?

OpenStudy (jojokiw3):

I can't really see what number one is, it's a bit bugged. Try reinputing it?

OpenStudy (study_buddy99):

\[\log_\frac{ 2 }{ 3 }\frac{ 27 }{ 8 }\]

OpenStudy (jojokiw3):

Ah. Yes, you're right.

OpenStudy (study_buddy99):

okay, so I get that concept just fine :) the problem is always the next part

OpenStudy (jojokiw3):

Will this help you? \[\frac{ a }{ b }^{-1} = \frac{ b }{ a }\]

OpenStudy (study_buddy99):

so in this case 2/1?

OpenStudy (jojokiw3):

\[\frac{ a }{ b }^{-x} = \frac{ b^x }{ a^x }\]

OpenStudy (study_buddy99):

-2/1

OpenStudy (jojokiw3):

Ok. So we have \[\frac{ 2 }{ 3 }^x = \frac{ 27 }{ 8 }\] We have to find one power that both 2 and 3 can be raised to in order to get 27 and 8. \[3^3 = 27, 2^3 = 8\] Now if you want to flip it, you merely have to add a negative sign to the exponent.

OpenStudy (study_buddy99):

so make them both ^-3?

OpenStudy (jojokiw3):

\[\frac{ 2 }{ 3 }^3 = \frac{ 8 }{ 27 }\] So now you just need to get \[\frac{ 27 }{ 8 }\] which can be done by changing the exponent to a negative.

OpenStudy (study_buddy99):

wait so the answer is -3?

OpenStudy (jojokiw3):

Yes \[\frac{ 2 }{ 3 }^ {-3} = \frac{ 27 }{ }\]

OpenStudy (jojokiw3):

the 8 is missing lol. o_O but yeah.

OpenStudy (study_buddy99):

ah okay! So for a more simple problem such as log_8 4 I know the answer is 2/3 but I'm not sure where I got the answer is it because 2^3 is 8 so then you make the answer 2-3?

OpenStudy (study_buddy99):

2/3*

OpenStudy (jojokiw3):

\[8^\frac{ 1 }{ 3 } = \sqrt[3]{8} = 2\]

OpenStudy (study_buddy99):

okay I see! So is number 2 any different?

OpenStudy (jojokiw3):

Different from what? o_O So \[8^\frac{ 2 }{ 3 }= \sqrt[3]{8} * \sqrt[3]{8} = 2*2 = 4\]

OpenStudy (study_buddy99):

I meant problem 2 :D

OpenStudy (jojokiw3):

Well no it's not, other than the negative. This may help \[x^{-\frac{ 1 }{ y }} = \frac{ 1 }{ \sqrt[y]{x} }\]and \[x^\frac{ 1 }{ y } = \sqrt[y]{x}\]

OpenStudy (jojokiw3):

So the numerator works like an exponent, but the denominator works as the "y" root of x. Does that make sense?

OpenStudy (jojokiw3):

Problem 2 is basically \[\frac{ 1 }{ 9 } = x^{-\frac{ 1 }{ 2 }}\]

OpenStudy (study_buddy99):

yes it does. so is it set up as: \[\sqrt[9]{x}=\frac{ 1 }{ \sqrt{x} }\]

OpenStudy (jojokiw3):

woah woah where did \[\sqrt[9]{x} \] come from?

OpenStudy (jojokiw3):

Now \[\sqrt[9]{x} = x^\frac{ 1 }{ 9 }\] but \[\sqrt[9]{x} \neq \frac{ 1 }{ 9 }\]

OpenStudy (study_buddy99):

so what do I do with x in this problem?

OpenStudy (jojokiw3):

\[\frac{ 1 }{ 9 } = \frac{ 1 }{ \sqrt{x} }\]

OpenStudy (study_buddy99):

so is the final answer 3?

OpenStudy (jojokiw3):

:P Check again.

OpenStudy (jojokiw3):

That's a square root, not a square, hehe.

OpenStudy (study_buddy99):

so it would be 81 :D

OpenStudy (jojokiw3):

Yeah haha.

OpenStudy (study_buddy99):

.-. I'm so tired... I can't math... but yet I need to do 50 more problems...

OpenStudy (jojokiw3):

hehe same. good luck!

OpenStudy (study_buddy99):

thank you, I need all that I can get. Can you explain the 3rd problem? I think the answer is 1/2...?